# Models That Prove Their Own Correctness

### v3

@@ Abstract
How can we trust the correctness of a learned model on a particular input of interest? Model accuracy is typically measured on average over a distribution of inputs, giving no guarantee for any fixed input. This paper proposes a theoretically-founded solution to this problem: to train Self-Proving models that prove the correctness of their output to a verification algorithm $V$ via an Interactive Proof.
Self-Proving models satisfy that, with high probability over an input sampled from a given distribution, the model generates a correct output and successfully proves its correctness to $V\!$. The soundness property of $V$ guarantees that, for every input, no model can convince $V$ of the correctness of an incorrect output. Thus, a Self-Proving model proves correctness of most of its outputs, while all incorrect outputs (of any model) are detected by $V$. We devise a generic methods for learning Self-Proving models, and prove its convergence under certain assumptions.
The theoretical framework and results are complemented by experiments on an arithmetic capability: computing the greatest common divisor (GCD) of two integers. Our learning method is used to train a Self-Proving transformer that computes the GCD and proves the correctness of its answer. Our code is available at https://github.com/orrp/self-proving-models.
@@ 1 Introduction
Bob is studying for his algebra exam and stumbles upon a question $Q$ that he cannot solve. He queries a Large Language Model (LLM) for the answer, and it responds with a number: 42. Bob is aware of recent research showing that the LLM attains a 90% score on algebra benchmarks (cf. Frieder et al. 2023), but should he trust that the answer to his particular question $Q$ is indeed 42?
Bob could ask the LLM to explain its answer in natural language. Though he must proceed with caution, as the LLM might try to convince him of an incorrect answer (Turpin et al., 2023). Moreover, even if 42 is the correct answer, the LLM may fail to produce a convincing proof (Wang et al., 2023). If only the LLM could formally prove its answer, Bob would verify the proof and be convinced.
This paper initiates the study of Self-Proving models (Fig. 1) that prove the correctness of their answers via an Interactive Proof system (Goldwasser et al., 1985). Self-Proving models successfully convince a verification algorithm $V$ with worst-case soundness guarantees: for any question, $V$ rejects all incorrect answers with high probability over the interaction. This guarantee holds even against provers that have access to $V$’s specification, and unbounded computational power.
@@
Guarantee Type Def.
$V$ Completeness & Soundness Worst-case $\forall x,y$ 3.2
$P_{\theta}$ Verifiability Average-case $x\sim\mu$, $y\sim P_{\theta}(x)$ 3.4
@@
  • We define Self-Proving models (Section 3).
  • We propose two methods for learning Self-Proving models in Section 4. The first, Transcript Learning (TL), relies on access to transcripts of accepting interactions and is the focus of this paper; we prove convergence bounds for TL under convexity and Lipschitzness assumptions. The second method, Reinforcement Learning from Verifier Feedback (RLVF), trains a model by emulating interaction with the verifier. We also present variants of these algorithms that use Annotations to improve learning in practice.
  • We empirically study TL and Annotated-TL (ATL) for training Self-Proving transformers that compute the Greatest Common Divisor (GCD) of two integers. Table 2 demonstrates the efficacy of our methods, with additional experiments in Section 5. Our results may be of independent interest for research on the arithmetic capabilities of transformers (e.g. Charton 2024; Lee et al. 2024). Code, data and models are available at https://github.com/orrp/self-proving-models.
Scope.
This paper contains a theory of learned models that prove their own correctness via an Interactive Proof system. The fascinating and well-studied question of which settings are verifiable in an Interactive Proof system is beyond our scope. Our theory is general in that it pertains to any such setting, e.g., any decision problem solvable in polynomial space (Shamir, 1992). See Goldreich (2008) for a primer on Proof systems more broadly.
Learning method Correctness (%) Verifiability (%)
GPT (baseline) 99.8 -
GPT+TL 98.8 60.3
GPT+TL+RLVF 98.9 78.3
GPT+Annotated TL 98.6 96.0
Table 2: Self-Proving transformers computing the GCD. We train a 6.3M parameter GPT to compute the GCD of two integers sampled log-uniformly from $[10^{4}]$. Vanilla GPT correctly generates the GCD for almost all inputs, but does not prove correctness to a simple verification algorithm. GPT trained with Transcript Learning (GPT+TL) proves its answer 60.3% of the time; adding Reinforcement Learning from Verifier Feedback (+RLVF) increases this to 78.3%; training with Annotated Transcript Learning (GPT+ATL) gives the highest Verifiability score of 96%. See Section 5 for details.
@@ 2 Related Work
This paper is situated at the intersection of machine learning (ML) and Interactive Proof systems (IPs). We briefly discuss recent relevant work from these literatures.
@@
IPs have found numerous applications in ML towards a diverse set of goals. Anil et al. (2021) introduce Prover–Verifier Games (PVGs), a game-theoretic framework for learned provers and learned verifiers. Since our paper initially appeared, PVGs were further investigated in at least two subsequent works: Hammond and Adam-Day (2024) study multi-prover and Zero Knowledge variants of PVGs. Additionally, Kirchner et al. (2024) successfully utilize PVGs towards obtaining human-legible outputs from LLMs. Notably, they require a relaxed completeness guarantee of their learned proof system—this requirement is the same as our Definition 3.4 of Self-Proving models.
Beyond PVGs, Wäldchen et al. (2024) cast the problem of model interpretability as a Prover–Verifier interaction between a learned feature selector and a learned feature classifier. Debate systems (Condon et al., 1995), a multiprover variant of IPs, were considered for aligning models with human values (Irving et al., 2018; Brown-Cohen et al., 2023). In such Debate systems, two competing models are each given an alleged answer $y\neq y^{\prime}$, and attempt to prove the correctness of their answer to a (human or learned) judge. Lastly, Murty et al. (2023) define Pseudointelligence: a model learner $L_{M}$ and an evaluator learner $L_{E}$ are each given samples from a ground-truth; $L_{M}$ learns a model of the ground-truth, while $L_{E}$ learns an evaluator of such models; the learned evaluator then attempts to distinguish between the learned model and the ground-truth in a Turing Test-like interaction.
All of these works consider learned verifiers, whereas our work focuses on training models that interact with a manually-defined verifier. More related in this regard is IP-PAC (Goldwasser et al., 2021), in which a learner proves that she learned a model that is Probably Approximately Correct (Valiant, 1984). We, however, consider models that prove their own correctness on a per-input basis, rather than learners that prove average-case correctness of a model.
@@
Self-Proving models are verified by an algorithm with formal completeness and soundness guarantees (see Definition 3.2). In this sense, Self-Proving models generate a formal proof of the correctness of their output. Several works propose specialized models that generate formal proofs.
AlphaGeometry (Trinh et al., 2024) is capable of formally proving olympiad-level geometry problems; Others have trained models to produce proofs in Gransden et al. (2015); Polu and Sutskever (2020) and others train models to produce proofs in Coq (Gransden et al., 2015), Metamath (Polu and Sutskever, 2020), Lean (Yang et al., 2023), or manually-defined deduction rules (Tafjord et al., 2020); FunSearch (Romera-Paredes et al., 2024) evolves LLM-generated programs by systematically evaluating their correctness. Indeed, all of these can be cast as Self-Proving models developed for specific proof systems. Meanwhile, this work defines and studies the class of such models in general. Several works (e.g. Welleck et al. 2022) consider models that generate natural language proofs or explanations, which are fundamentally different from formal proofs (or provers) verified by an algorithm.
@@
Chain-of-Though (CoT, Wei et al. 2022) refers to additional supervision on a model in the form of intermediate reasoning steps. CoT is known to improve model performance whether included in-context (Wei et al., 2022) or in the training phase itself (Yang et al., 2022). Transcript Learning (TL, Section 4.1) can be viewed as training the model on a Chain-of-Thought induced by the interaction of a verifier and an honest prover (Definition 3.2).
To complete the analogy, let us adopt the terminology of Uesato et al. (2022), who consider outcome supervision and process supervision. In our case, the outcome is the decision of the verifier, and the process is the interaction between the verifier and the model. Thus, Reinforcement Learning from Verifier Feedback (RLVF, Section 4.2) is outcome-supervised while TL is process-supervised. In a recent work, Lightman et al. (2024) find that process-supervised transformers outperform outcome-supervised ones on the MATH dataset (Hendrycks et al., 2021).
Transformers for arithmetic.
In Section 5 we train and evaluate Self-Proving transformers to generate the GCD of two integers and prove its correctness to a verifier. These experiments leverage a long line of work on neural models for arithmetic tasks originating with Siu and Roychowdhury (1992), and in particular modular arithmetic, which is known to be challenging (Palamas, 2017). Of particular relevance is the recent paper of Charton (2024), who trains transformers to generate the GCD—without a proof of correctness. We benefit from conclusions suggested in their work and start from a similar (scaled-down) experimental setup. Our main challenge (obtaining Self-Proving models) is overcome by introducing Annotated Transcript Learning (ATL).
We conduct ablation experiments to find two deciding factors in ATL. First, we study the effect of the amount of annotation given in the form of intermediate steps (Lee et al., 2024), which is related to autoregressive length complexity (Malach, 2023). Second, we characterize ATL efficacy in terms of an algebraic property of the tokenization scheme (cf. Nogueira et al. 2021; Charton 2022, 2024).
3 Self-Proving models
We introduce and formally define our learning framework in which models prove the correctness of their output. We start with preliminaries from the learning theory and proof systems literatures in Section 3.1. We then introduce our main definition in Section 3.2.
@@ 3.1 Preliminaries
Before we can define models that prove their own correctness, we must first define correctness. Correctness is defined with respect to an input distribution $\mu$ over $\Sigma^{\ast}$, and a ground-truth $F^{\ast}$ that defines correct answers. For simplicity of presentation, we focus on the case that each input $x\in\Sigma^{\ast}$ has exactly one correct output $F^{\ast}(x)\in\Sigma^{\ast}$, and a zero-one loss function on outputs (the general case is deferred to Appendix A). The fundamental goal of machine learning can be thought of as learning a model of the ground-truth $F^{\ast}$. Formally,
@@
An interactive proof system (Goldwasser et al., 1985) is a protocol carried out between an efficient verifier and a computationally unbounded prover. The prover attempts to convince the verifier of the correctness of some assertion, while the verifier accepts only correct claims. The prover is powerful yet untrusted; in spite of this, the verifier must reject false claims with high probability.
@@
A Self-Proving model proves that $y\sim P_{\theta}(x)$ is correct to a verifier $V$ over the course of $R$ rounds of interaction (Figure 1). In each round $i\in[R]$, verifier $V$ queries $P_{\theta}$ on a sequence $q_{i}\in\Sigma^{\ast}$ to obtain an answer $a_{i}\in\Sigma^{\ast}$; once the interaction is over, $V$ accepts or rejects. For fixed $x,y\in\Sigma^{\ast}$, the decision of $V$ after interacting with $P_{\theta}$ is a random variable over $V$’s decision (accept/reject), determined by the randomness of $V$ and $P_{\theta}$. The decision random variable is denoted by $V,P_{\theta}\right>(x,y)$.
We present a definition of Interactive Proofs restricted to our setting.
@@
Fix a soundness error $s\in(0,1)$, a finite set of tokens $\Sigma$ and a ground-truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. A verifier $V$ (in an Interactive Proof) for $F^{\ast}$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover $P$ [^2] We intentionally write $P$ rather than $P_{\theta}$: Interactive Proofs are defined with respect to all possible provers, not just parameterized ones. It interacts with $P$ over $R$ rounds (see Figure 1) and outputs a decision $V,P\right>(x,y)\in\{\mathrm{reject},\mathrm{accept}\}$. Verifier $V$ satisfies the following two guarantees:
@@
    where the probability is over the randomness of $V$ [^3] WLOG, the honest prover is deterministic by fixing the optimal randomness of a randomized prover.
@@
Definition 3.2 requires that $V$ is a polynomial-time algorithm whereas provers are unbounded. This captures a requirement for efficient verification. We chose polynomial time as a measure of efficiency because it is common Proof systems literature. That said, one could adapt Definition 3.2 to fit alternative efficiency measures, such as space complexity (Condon and Lipton, 1989) or circuit depth (Goldwasser et al., 2007). Regardless of which measure is taken, to avoid a trivial definition it is crucial that $V$ should be more efficient than the honest prover $P^{\ast}$; else, $V$ can simply execute $P^{\ast}$ to perform the computation itself.
@@
As depicted in Figure 1, each of the model’s answers depends on all previous queries and answers in the interaction. This captures the setting of stateful models, e.g. a session with a chatbot.
Towards defining Self-Proving models (Section 3.2), let us observe the following. Completeness and soundness are worst-case guarantees, meaning that they hold for all possible inputs $x\in\Sigma^{\ast}$. In particular, completeness implies that for all $x\in\Sigma^{\ast}$, the honest prover $P^{\ast}$ convinces $V$ of the correctness of $F^{\ast}(x)$; in classical proof systems there is no guarantee that an “almost honest” prover can convince the verifier (cf. Paradise 2021). Yet, if we are to learn a prover $P_{\theta}$, we cannot expect it to agree with $P^{\ast}$ perfectly, nor can we expect it to always output $F^{\ast}(x)$. Indeed, Self-Proving models will have a distributional guarantee with respect to inputs $x\sim\mu$.

### v4

@@ Abstract
How can we trust the correctness of a learned model on a particular input of interest? Model accuracy is typically measured on average over a distribution of inputs, giving no guarantee for any fixed input. This paper proposes a theoretically-founded solution to this problem: to train Self-Proving models that prove the correctness of their output to a verification algorithm $V$ via an Interactive Proof. Self-Proving models satisfy that, with high probability over an input sampled from a given distribution, the model generates a correct output and successfully proves its correctness to VV\!. The soundness property of $V$ guarantees that, for every input, no model can convince $V$ of the correctness of an incorrect output. Thus, a Self-Proving model proves correctness of most of its outputs, while all incorrect outputs (of any model) are detected by $V$. We devise and analyze two generic methods for learning Self-Proving models: Transcript Learning (TL) which relies on access to transcripts of accepting interactions, and Reinforcement Learning from Verifier Feedback (RLVF) which trains a model by emulating interactions with the verifier.
@@ 1 Introduction
Bob is studying for his algebra exam and stumbles upon a question $Q$ that he cannot solve. He queries a Large Language Model (LLM) for the answer, and it responds with a number: 42. Bob is aware of recent research showing that the LLM attains a 90% score on algebra benchmarks (cf. FriederPCGSLPB23), but should he trust that the answer to his particular question $Q$ is indeed 42?
Bob could ask the LLM to explain its answer in natural language. Though he must proceed with caution, as the LLM might try to convince him of an incorrect answer (TurpinMPB23). Moreover, even if 42 is the correct answer, the LLM may fail to produce a convincing proof (WangY023). If only the LLM could formally prove its answer, Bob would verify the proof and be convinced.
This paper initiates the study of Self-Proving models (Fig.˜1) that prove the correctness of their answers via an Interactive Proof system (GoldwasserMR85). Self-Proving models successfully convince a verification algorithm $V$ with worst-case soundness guarantees: for any question, $V$ rejects all incorrect answers with high probability over the interaction. This guarantee holds even against provers that have access to $V$’s specification, and unbounded computational power.
@@
  • We define Self-Proving models (Section˜3).
  • We propose two methods for learning Self-Proving models (Section˜4). The first, Transcript Learning (TL), relies on access to transcripts of accepting interactions. The second method, Reinforcement Learning from Verifier Feedback (RLVF), trains a model by emulating interactions with the verifier.
  • We prove gradient approximation lemmas for both methods (Lemmas˜4.2 and 4.3), and a convergence bounds for TL under convexity and Lipschitzness assumptions (Section˜5). These are supplemented by empirical validation on a simple arithmetic capability (Section˜7). Code and data are at https://github.com/orrp/self-proving-models
This paper develops a theory of learned models that prove their own correctness via an interactive proof system, and thus lies at the intersection of machine learning and Interactive Proof systems. We defer the discussion of relevant prior work from these areas to the related work section in Section˜2. The rich and well-studied question of which settings are verifiable within an interactive proof system is beyond our scope. Our theory is general in that it applies to any setting which is verifiable within an interactive proof system, e.g., any decision problem solvable in polynomial space (Shamir92). For a broader introduction to proof systems, see Goldreich08primer.
@@ 2 Related Work
We overview related work from machine learning (ML) and interactive proof systems (IPs) literature.
@@
IPs have found numerous applications in ML towards a diverse set of goals. AnilZWG21 introduce Prover–Verifier Games (PVGs), a game-theoretic framework for learned provers and learned verifiers. Since our paper initially appeared, PVGs were further investigated in at least two subsequent works: HammondD24 study multi-prover and Zero Knowledge variants of PVGs. Additionally, KirchnerCELMB24 successfully utilize PVGs towards obtaining human-legible outputs from LLMs. Notably, they require a relaxed completeness guarantee of their learned proof system—this requirement is the same as our Definition˜3.4 of Self-Proving models.
Beyond PVGs, Waldchen24a cast the problem of model interpretability as a Prover–Verifier interaction between a learned feature selector and a learned feature classifier. Debate systems (CondonFLS95), a multiprover variant of IPs, were considered for aligning models with human values (IrvingCA18; BrownIP23). In such Debate systems, two competing models are each given an alleged answer $y\neq y^{\prime}$, and attempt to prove the correctness of their answer to a (human or learned) judge. Lastly, MurtyPS23 define Pseudointelligence: a model learner $L_{M}$ and an evaluator learner $L_{E}$ are each given samples from a ground-truth; $L_{M}$ learns a model of the ground-truth, while $L_{E}$ learns an evaluator of such models; the learned evaluator then attempts to distinguish between the learned model and the ground-truth in a Turing Test-like interaction.
All of these works consider learned verifiers, whereas our work focuses on training models that interact with a manually-defined verifier. More related in this regard is IP-PAC (GoldwasserRSY21), in which a learner proves that she learned a model that is Probably Approximately Correct (Valiant84). We, however, consider models that prove their own correctness on a per-input basis, rather than learners that prove average-case correctness of a model.
@@
Self-Proving models are verified by an algorithm with formal completeness and soundness guarantees (see Definition˜3.2). In this sense, Self-Proving models generate a formal proof of the correctness of their output. Several works propose specialized models that generate formal proofs.
AlphaGeometry (TrinhWLHL24) is capable of formally proving olympiad-level geometry problems; Others have trained models to produce proofs in GransdenWR15; PoluS20 and others train models to produce proofs in Coq (GransdenWR15), Metamath (PoluS20), Lean (YangSGCSYGPA23), or manually-defined deduction rules (TafjordDC20); FunSearch (Romera24) evolves LLM-generated programs by systematically evaluating their correctness. Indeed, all of these can be cast as Self-Proving models developed for specific proof systems. Meanwhile, this work defines and studies the class of such models in general. Several works (e.g. WelleckLLHC22) consider models that generate natural language proofs or explanations, which are fundamentally different from formal proofs (or provers) verified by an algorithm.
@@
Chain-of-Though (CoT, Wei0SBIXCLZ22) refers to additional supervision on a model in the form of intermediate reasoning steps. CoT is known to improve model performance whether included in-context (Wei0SBIXCLZ22) or in the training phase itself (YangSAN22). Transcript Learning (TL, Section˜4.1) can be viewed as training the model on a Chain-of-Thought induced by the interaction of a verifier and an honest prover (Definition˜3.2).
To complete the analogy, let us adopt the terminology of Uesato22, who consider outcome supervision and process supervision. In our case, the outcome is the decision of the verifier, and the process is the interaction between the verifier and the model. Thus, Reinforcement Learning from Verifier Feedback (RLVF, Section˜4.2) is outcome-supervised while TL is process-supervised. In a recent work, LightmanKBEBLLSSC24 find that process-supervised transformers outperform outcome-supervised ones on the MATH dataset (HendrycksBKABTS21).
Subsequent work on RLVF.
Following the preprint version of this paper on May 2024, Reinforcement Learning from Verifier Feedback (RLVF, Section˜4.2) has been implemented and widely adopted. Of particular note is RLVR (Reinforcement Learning from Verifiable Reward, LambertEtAl24), which implements RLVF by adding KL regularization to the RL objective of Algorithm˜2. This implementation provided valuable empirical validation and state-of-the-art performance on full-scale LLMs, sparked additional empirical analysis (WuXLHC25; zhou2025rl), and led to widespread adoption across numerous domains including medical reasoning, computer vision, and robotics (med-rlvf; vision-rlvf; robotics-rlvf). Most recently, PyatkinEtAl25 presented IFBench which measures performance in interactive (3-round) proof systems. These empirical contributions complement our theoretical foundations, demonstrating the surprising power and applicability of RLVF as a practical post-training method.
3 Defining Self-Proving Models
We introduce and formally define our learning framework in which models prove the correctness of their output. We start with preliminaries from the learning theory and proof systems literatures in Section˜3.1. We then introduce our main definition in Section˜3.2.
@@ 3.1 Preliminaries
Before we can define models that prove their own correctness, we must first define correctness. Correctness is defined with respect to an input distribution $\mu$ over $\Sigma^{\ast}$, and a ground-truth $F^{\ast}$ that defines correct answers. For simplicity of presentation, we focus on the case that each input $x\in\Sigma^{\ast}$ has exactly one correct output $F^{\ast}(x)\in\Sigma^{\ast}$, and a zero-one loss function on outputs (the general case is deferred to Section˜3.3). The fundamental goal of machine learning can be thought of as learning a model of the ground-truth $F^{\ast}$. Formally,
@@
An interactive proof system (GoldwasserMR85) is a protocol carried out between an efficient verifier and a computationally unbounded prover. The prover attempts to convince the verifier of the correctness of some assertion, while the verifier accepts only correct claims. The prover is powerful yet untrusted; in spite of this, the verifier must reject false claims with high probability.
@@
A Self-Proving model proves that $y\sim P_{\theta}(x)$ is correct to a verifier $V$ over the course of $R$ rounds of interaction (Figure˜1). In each round $i\in[R]$, verifier $V$ queries $P_{\theta}$ on a sequence $q_{i}\in\Sigma^{\ast}$ to obtain an answer $a_{i}\in\Sigma^{\ast}$; once the interaction is over, $V$ accepts or rejects. For fixed $x,y\in\Sigma^{\ast}$, the decision of $V$ after interacting with $P_{\theta}$ is a random variable over $V$’s decision (accept/reject), determined by the randomness of $V$ and $P_{\theta}$. The decision random variable is denoted by $\left<V,P_{\theta}\right>(x,y)$.
Next, we present a definition of Interactive Proofs restricted to our setting.
@@
Fix a soundness error $s\in(0,1)$, a finite set of tokens $\Sigma$ and a ground-truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. A verifier $V$ (in an Interactive Proof) for $F^{\ast}$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover PP. [^2] We intentionally write $P$ rather than $P_{\theta}$: Interactive Proofs are defined with respect to all possible provers, not just parameterized ones. It interacts with $P$ over $R$ rounds (see Figure˜1) and outputs a decision $\left<V,P\right>(x,y)\in\{\mathrm{reject},\mathrm{accept}\}$. Verifier $V$ satisfies the following two guarantees:
@@
    where the probability is over the randomness of VV. [^3] WLOG, the honest prover is deterministic by fixing the optimal randomness of a randomized prover.
@@
Definition˜3.2 requires that $V$ is a polynomial-time algorithm whereas provers are unbounded. This captures a requirement for efficient verification. We chose polynomial time as a measure of efficiency because it is common in the Proof systems literature. That said, one could adapt Definition˜3.2 to fit alternative efficiency measures, such as space complexity (CondonL89) or circuit depth (GoldwasserGHKR07). Regardless of which measure is taken, to avoid a trivial definition it is crucial that $V$ should be more efficient than the honest prover $P^{\ast}$; else, $V$ can simply execute $P^{\ast}$ to perform the computation itself.
@@
As depicted in Figure˜1, each of the model’s answers depends on all previous queries and answers in the interaction. This captures the setting of stateful models, e.g. a session with a chatbot.
Towards defining Self-Proving models (Section˜3.2), let us observe the following. Completeness and soundness are worst-case guarantees, meaning that they hold for all possible inputs $x\in\Sigma^{\ast}$. In particular, completeness implies that for all $x\in\Sigma^{\ast}$, the honest prover $P^{\ast}$ convinces $V$ of the correctness of $F^{\ast}(x)$; in classical proof systems there is no guarantee that an “almost honest” prover can convince the verifier (cf. Paradise21). Yet, if we are to learn a prover $P_{\theta}$, we cannot expect it to agree with $P^{\ast}$ perfectly, nor can we expect it to always output $F^{\ast}(x)$. Indeed, Self-Proving models will have a distributional guarantee with respect to inputs $x\sim\mu$. This distinction is summarized in Table˜1.
Guarantee Type Def.
$V$ Completeness & Soundness Worst-case: $\forall x,y$ 3.2
$P_{\theta}$ Verifiability Average-case: $x\sim\mu,\ y\sim P_{\theta}(x)$ 3.4